# thanks: These authors contributed equallythanks: These authors contributed equally

### v1

@@ Abstract
Fast, reliable logical operations are essential for the realization of useful quantum computers [1, 2, 3], as they are required to implement practical quantum algorithms at large scale. By redundantly encoding logical qubits into many physical qubits and using syndrome measurements to detect and subsequently correct errors, one can achieve very low logical error rates. However, for most practical quantum error correcting (QEC) codes such as the surface code, it is generally believed that due to syndrome extraction errors, multiple extraction rounds—on the order of the code distance $d$—are required for fault-tolerant computation [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. Here, we show that contrary to this common belief, fault-tolerant logical operations can be performed with constant time overhead for a broad class of QEC codes, including the surface code with magic state inputs and feed-forward operations, to achieve “algorithmic fault tolerance”. Through the combination of transversal operations [7] and novel strategies for correlated decoding [15], despite only having access to partial syndrome information, we prove that the deviation from the ideal measurement result distribution can be made exponentially small in the code distance. We supplement this proof with circuit-level simulations in a range of relevant settings, demonstrating the fault tolerance and competitive performance of our approach. Our work sheds new light on the theory of quantum fault tolerance, potentially reducing the space-time cost of practical fault-tolerant quantum computation by orders of magnitude.
@@
Performing large-scale computation, however, comes with significant overhead [16, 2]. By performing syndrome extraction (SE), one can reveal error information and use a classical decoder to correct physical errors in software and interpret logical measurement results. However, in the presence of noisy syndrome measurements [4, 5, 6, 7, 10], one typically requires a number of SE rounds that scales linearly in $d$, i.e., $\Theta(d)$  [^1] The notation $g(x)=\Theta(f(x))$ indicates that two functions $f(x)$ and $g(x)$ have the same asymptotic scaling with $x$, or more precisely, that there exists some constants $c_{1}$ and $c_{2}$ such that $c_{1}f(x)\leq g(x)\leq c_{2}f(x)$ for sufficiently large $x$. (see Fig. 1(a)). This is the case, for example, for the celebrated surface code [8, 9, 10], one of the leading candidates for practical FTQC due to its simple 2D layout and competitive error thresholds. In typical compilations based on lattice surgery or braiding [11, 12, 13, 14, 18], each logical operation requires $\Theta(d)$ SE rounds, thus incurring a space-time volume per logical operation of $\Theta(d^{3})$. This reduces the logical clock speed by a factor proportional to the code distance, typically on the order of 10 –100 [16, 14]. The same considerations also apply when performing logical operations with many quantum low-density parity-check (QLDPC) codes [19, 20]. While there have been various efforts at addressing this challenge [5, 21], these alternative approaches introduce higher hardware complexity [22, 23, 20, 24] or necessitate certain properties of the underlying codes, such as the single shot QEC property [25, 26, 27, 28, 29], often incurring a trade-off between space and time when executing logical operations [30, 16, 2].
@@
Figure 1: Algorithmic fault tolerance. (a) Conventional FT analysis separately examines each gadget (red boxes) in the circuit and ensures they are individually FT [4, 7, 31]. This requires $\Theta(d)$ syndrome extraction (SE) rounds to achieve FT. (b) Algorithmic FT directly uses all accessible syndrome information up to a logical measurement (blue box), and guarantees FT of the measurement result, even if the gadgets are not individually FT and if future syndrome information is not yet accessible (partial decoding). We realize algorithmic FT through transversal operations, and only require a single SE round per logical operation, thus allowing constant time implementations of logical operations.
@@
We introduce and develop a novel approach to FTQC that we refer to as “algorithmic fault tolerance”, and show that it can lead to a substantial reduction in space-time cost. We focus on transversal implementations of Clifford circuits [7, 32] with magic state inputs and feed-forward [33], thereby allowing universal quantum computation. Such transversal gate capabilities have already been demonstrated in multiple hardware platforms, such as neutral atoms and trapped ions [34, 35, 36]. We show that contrary to the common belief, for any Calderbank-Shor-Steane (CSS) QLDPC code [6, 37], these operations can be performed fault-tolerantly with only constant time overhead per operation, provided that decoding can be implemented efficiently. The key idea is to consider the fault tolerance of the algorithm as a whole (Fig. 1(b)) [38, 39, 40]. We achieve this by performing correlated decoding [15, 36, 30] despite only having access to partial syndrome information, and ensuring consistency in the presence of magic states and feed-forward via additional operations in software. We verify such algorithmic fault tolerance through a combination of proofs and circuit-level numerical simulations of our protocol, including a simulation of state distillation factories [33, 13], finding very little change to physical error thresholds. Specializing to the surface code, our results reduce the per-operation time cost from $\Theta(d)$ to $\Theta(1)$, including for Clifford operations used in magic state distillation. Note that unlike methods that trade space for time, our techniques represent a direct reduction in space-time volume, which is usually the ultimate quantity of interest.
Algorithmic Fault Tolerance via Transversal Operations
We focus on transversal Clifford circuits with magic state inputs, where Clifford operations are implemented with a depth-one quantum circuit (Methods). This is interleaved with SE rounds using ancilla qubits, which reveal error information on the data qubits and enable error correction. In addition to transversal gates [7], we refer to preparation of data qubits in $|0\rangle$ followed by one SE round as transversal state preparation, and $Z$ basis measurement of all data qubits as transversal measurement. To achieve universality, we allow teleporting in low-noise magic states with feed-forward operations based on past measurement results, and use the same Clifford operations above to prepare high quality magic states via magic state distillation [33]. We make use of CSS QLDPC codes, where each data or ancilla qubit interacts with a constant number of other qubits, and each stabilizer generator consists of all $X$ or all $Z$ operators [37, 6]. Within this setting, our key result can be formulated as the following theorem:
Theorem 1 (informal): Exponential error suppression for constant time transversal Clifford operations with any CSS QLDPC code. For a transversal Clifford circuit with low-noise magic state inputs and feed-forward operations, that can be implemented with a given CSS QLDPC code family $\mathcal{Q}_{d}$ of growing code distance $d$, there exists a threshold $p_{\mathrm{th}}$, such that if the physical error rate $p<p_{\mathrm{th}}$ under the basic model of fault tolerance [5], then our protocol can perform constant time logical operations, with only a single SE round per operation, while suppressing the total logical error rate as $P_{L}=\exp(-\Theta(d_{n}))$.
The formal theorem statement and the corresponding proof can be found in Supplementary Materials [41]. Our analysis assumes the basic model of fault tolerance [5]. In particular, we consider the local stochastic noise model, where we apply depolarizing errors on each data qubit every SE round and measurement errors on each SE result, with a probability that decays exponentially in the weight of the error event. This can be readily generalized to circuit-level noise by noting the bounded error propagation for constant depth SE circuits in QLDPC codes. We also assume the most likely error (MLE) decoder and fast classical computation (Methods). Finally, we assume that all code patches are identical, and the number of qubit locations within a code patch that any given qubit can be coupled to via transversal gates is bounded by some constant $t$, in order to control error propagation.
A key observation is that by considering the algorithm as a whole and leveraging the deterministic propagation of errors through transversal Clifford circuits, one can use the surrounding syndrome history to correct for noisy measurements (Fig. 1(b)). This correlated decoding technique has been shown to enable $\Theta(1)$ SE rounds for Clifford circuits without feed-forward [15]. However, a key component of many schemes for achieving universality is magic state teleportation, which crucially relies on the ability to realize feed-forward operations.
As illustrated by the example shown in Fig. 2(a), such feed-forward operations require on-the-fly interpretation of logical measurements, followed by a subsequent conditional gate, when only a subset of the logical qubits have been measured. As we do not yet have future syndrome information on the unmeasured logical qubits, one may be concerned that this can lead to an incorrect assignment of logical measurement results. Indeed, prior work analyzing circuits with magic states assumed that at least $d$ SE rounds separated state initialization and measurements or out-going qubits [30, 42, 43]. As shown in Fig. 2(b) for the $\Theta(1)$ SE round case, with new syndrome information, one may end up concluding a different measurement result, which leads to an incorrect feed-forward operation.
Surprisingly, we find that these inconsistencies can be accounted for in classical processing, with a reinterpretation of subsequent measurement results (Fig. 2(c), Pauli frame updates). The inconsistent measurement result corresponds to an $\overline{X}$ operator applied right before the $\overline{Z}$ measurement. Tracing back, we can find an $\overline{X}$ operator on the $|{\overline{+}}\rangle$ initial state (Fig. 2(c)) which does not change the logical state but propagates through to apply $\overline{X}$ on the logical measurement, together with some other logical Pauli updates on the remaining logical qubits. These are stabilizers of the logical state, which leave the state invariant. Indeed, the fact that this measurement result can be affected by non-fault-tolerant state preparation implies that the measurement anti-commutes with the corresponding Pauli stabilizer, necessarily leading to a 50/50 random outcome that is not changed by a logical flip. Products of individual measurements can have nontrivial correlations only if they commute with all the Pauli stabilizers. Because they commute, however, they are also guaranteed to be insensitive to the state initialization errors.
Therefore, in the second step of our decoding procedure, we apply such Pauli operators on initial input states until the measurement results are consistent with the previous commitments (Fig. 2(c)). Beyond this specific circuit, the required pattern that leads to a consistent assignment can always be computed efficiently by solving a linear system of equations (Methods). In practice, subsets of measurements in which all measurement products are 50/50 random can be classically assigned in advance, with the future measurements determined through the above procedure to ensure consistency. This also implies that decoding of certain measurements can be delayed until joint products need to be determined, and some assignments can be performed deterministically in specific cases such as state distillation (Methods).
Our protocol that leads to Theorem 1 thus consists of two main steps: correlated decoding based on partial syndrome information, and application of logical stabilizers to guarantee consistency between multiple decoding rounds (Fig. 2).
@@
Figure 2: Illustration of decoding strategy. (a) Logical quantum circuit with measurement and feed-forward. All logical operations are transversal and interleaved with a single SE round, instead of $d$ SE rounds. We must decode and commit mid-circuit to a measurement result for the bottom qubit, despite lacking complete syndrome information on the top two qubits (partial decoding). (b) With the measurement result of the bottom qubit, a feed-forward operation is applied, the remaining circuit is executed, and decoding is performed again on the whole circuit. The second decoding round may assign a different result to the bottom qubit, causing an inconsistency in feed-forward operations. (c) To guarantee consistency, we apply an $\overline{X}$ operator on the $|{\overline{+}}\rangle$ initial state of the middle qubit, which acts trivially on $|{\overline{+}}\rangle$, but changes the interpreted logical measurement result $M_{a}$ to be consistent with before. This also leads to a re-interpretation of the logical measurement result $M_{2}$.
We now sketch the intuition behind our proof of Theorem 1. There are two types of logical errors that may occur with our protocol. The first, a heralded inconsistency error, occurs when we are not able to find a set of operators to apply that yield the same outcome as previously committed measurement results. The second, a regular error, occurs when an erroneous logical operator is applied that results in a different measurement distribution.
Because imperfect readout during transversal measurements are equivalent to data qubit errors followed by perfect measurements, transversal measurements produce reliable syndrome information. Intuitively, this prevents individual errors from leading to high-weight corrections on the logical qubits we measure, the main reason for needing $d$ SE rounds in typical FT state initialization protocols. At the same time, the use of correlated decoding, together with the structured error propagation through transversal Clifford gates, allow us to propagate this syndrome information and correct relevant errors happening throughout the circuit. With these observations, we prove that for either type of logical error to occur, the total Pauli weight $s$ of physical error and subsequent correction in a connected cluster must satisfy $s=\Theta(d)$, which has probability $p^{s/2}$ under the MLE decoder. Finally, we count the number of such connected clusters of size $s$, which scales as $(v\mathfrak{e})^{s}$, where $\mathfrak{e}$ is the natural base and $v$ is a constant upper-bounding the error connectivity for a QLDPC code. The combined probability of an error thus scales as
$P_{\mathrm{err}}\propto p^{s/2}(2v\mathfrak{e})^{s}=\quantity(2v\mathfrak{e}\sqrt{p})^{\Theta(d)}\rightarrow 0$ (1)
when the physical error rate is sufficiently low $p\ll 1/(v\mathfrak{e})^{2}$ (the factor of 2 comes from a combinatorial sum), thereby establishing the existence of a threshold and exponential error suppression.
@@
Specializing to the surface code and utilizing the full transversal Clifford gate set accessible to the surface code (Methods), an immediate corollary of our main theorem is a threshold result for performing constant time logical operations with an arbitrary transversal Clifford circuit. This result supports universal quantum computing when we allow magic state inputs prepared with sufficiently low noise.
Preparing high quality magic state inputs, in turn, can be performed simply with the same Clifford operations and easy-to-prepare non-fault-tolerant magic states [44, 45, 46], a procedure known as magic state distillation [33] (see ED Fig. 3). We expect that the same algorithmic FT approach described above achieves a $\Theta(d)$ speed-up in distillation time as well. The distillation factory and main computation can then be combined by applying our decoding approach to the joint system. In Methods and Supplementary Information, we further describe an extension of our results to the case of single-shot code patch growth, relevant to practical distillation factories [47, 48]. Taken together, these results provide a theoretical foundation for our factor of $\Theta(d)$ improvement in logical clock speed compared to standard FT approaches for universal quantum computation.
@@
Figure 3: Numerical verification of fault tolerance. (a) Simulation of circuit with repeated $\overline{ZZ}$ measurement (inset), where we commit mid-circuit to each measurement result of the logical ancilla using only the syndrome information up to that point. The total logical error rate as a function of circuit-level physical error rate $p$, for varying code distance $d$, shows clear threshold behavior. (b) Heralded error rate with and without the second step of our decoding strategy, as a function of code distance and for different physical error rates, for the same circuit as (a). Only with both steps do we observe exponential suppression of the logical error rate. (c) Comparison of two different methods for logical state preparation between three rotated surface codes and subsequent teleportation, for fixed circuit noise $p=0.3\%$. We use transversal gates (left) and lattice surgery (right), in both cases with only a single SE round. (d) With transversal gates, the error rate decreases exponentially with the code distance. With a single round of lattice surgery, the error rate instead increases linearly with code distance, as a single stabilizer measurement error affects the logical $\overline{ZZ}$ measurement result.
@@
Figure 4: $|{\overline{Y}}\rangle$ state distillation factory. (a) Illustration of a $|{\overline{Y}}\rangle$ state distillation factory based on the [[7,1,3]] Steane code, consisting of state initialization, layers of transversal CNOTs, followed by a teleported $\overline{S}$ gate. Each operation involves only a single SE round. Two of the CNOTs in the first layer act trivially and can be omitted. (b) The $|{\overline{Y}}\rangle$ resource state is prepared via state injection at the first level, and via the first-level factory for the second level. (c) 1-level factory output state infidelity as a function of input state infidelity, for fixed circuit noise $p=0.1\%$ and varying levels of artificially injected $Z$ errors. The ideal curve is calculated assuming the gate operations in the factory are perfect. (d) Performance for one and two rounds of distillation, showing good agreement with the expected scaling.
@@
We first consider a simple circuit with intermediate logical measurements (inset of Fig. 3(a)). In this example, two logical qubits are transversally initialized in $|{\overline{+}}\rangle$, and an ancilla logical qubit is used to measure the $\overline{Z}\overline{Z}$ correlation a total of eight times, before the two logical qubits are transversally measured in the $\overline{Z}$ basis. While individual logical measurement results are random, a correct realization of this circuit should yield the same result for $\overline{Z}\overline{Z}$ each time, which in turn should be consistent with the final logical measurement results. We employ our algorithmic FT protocol to decode the circuit up to each logical measurement using only the syndrome information accessible at that point. We use the rotated surface code, a circuit-level depolarizing noise model [49, 15], a MLE decoder based on integer programming [50, 15], and employ the two-step process described above (see Supplementary Information).
@@
Figure 3(a-b) show the results of numerical simulations. We find that the total logical error rate, defined as the probability that a logical error of either type mentioned above happened anywhere in the circuit, shows characteristic threshold behavior, with an estimated threshold $\gtrsim 0.85\%$. As an SE round involves four layers of CNOT gates, while the transversal CNOT only involves a single layer, the effective error rate is dominated by SE operations, hence it may be expected that the threshold is close to the circuit-noise memory threshold. The number of SE rounds can be further optimized: for example, in Ref. [15], performing one SE round every four gate layers minimized the space-time cost per CNOT, suggesting that the practical improvement may be $\gtrsim 2d$ in some regimes  [^2] Four transversal gates and one SE round require 8 CNOT gates in total, compared to $416d+$ 4 with $d$ SE rounds.. In Fig. 3(b), we further compare the scaling of heralded failure rates in the presence and absence of the second step of our decoding procedure, as a function of code distance $d$. We find that this additional step is crucial to achieve exponential suppression with the code distance.
@@
We now contrast our approach with lattice surgery in a similar setting [11, 12, 52, 18]. We consider the logical circuit in Fig. 3(c), where a GHZ state preparation circuit is followed by teleportation of the GHZ state to another set of logical qubits, and then measurement in the $\overline{Z}$ basis [41]. Using transversal gates with only a single SE round during $|{\overline{+}}\rangle$ and $|{\overline{0}}\rangle$ state preparation, and decoding each logical measurement with only accessible information at that stage, we find that the logical error rate decreases exponentially with the code distance, consistent with our FT analysis. In contrast, state preparation based on a single round of lattice surgery [52], which involves performing syndrome extraction with a larger code patch and then splitting it into three individual logical qubits, does not yield improved logical error rate as the code distance increases, as a single error can lead to incorrect inference of the $\overline{Z}\overline{Z}$ correlation of the GHZ state (Supplementary Information). Unlike transversal measurements, logical information here is contained in noisy stabilizer products, which require repetition to reliably infer.

### v2

@@ Abstract
Fast, reliable logical operations are essential for realizing useful quantum computers [1, 2, 3]. By redundantly encoding logical qubits into many physical qubits and using syndrome measurements to detect and correct errors, one can achieve low logical error rates. However, for many practical quantum error correcting (QEC) codes such as the surface code, due to syndrome measurement errors, standard constructions require multiple extraction rounds—on the order of the code distance dd—for fault-tolerant computation, particularly considering fault-tolerant state preparation [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. Here, we show that logical operations can be performed fault-tolerantly with only a constant number of extraction rounds for a broad class of QEC codes, including the surface code with magic state inputs and feed-forward, to achieve “transversal algorithmic fault tolerance”. Through the combination of transversal operations [7] and novel strategies for correlated decoding [15], despite only having access to partial syndrome information, we prove that the deviation from the ideal logical measurement distribution can be made exponentially small in the distance, even if the instantaneous quantum state cannot be made close to a logical codeword due to measurement errors. We supplement this proof with circuit-level simulations in a range of relevant settings, demonstrating the fault tolerance and competitive performance of our approach. Our work sheds new light on the theory of quantum fault tolerance and has the potential to reduce the space-time cost of practical fault-tolerant quantum computation by over an order of magnitude.
@@
Performing large-scale computation, however, comes with significant overhead [16, 2]. By performing syndrome extraction (SE), one can reveal error information and use a classical decoder to correct physical errors in software and interpret logical measurement results. However, in the presence of noisy syndrome measurements [4, 5, 6, 7, 10], one typically requires a number of SE rounds that scales linearly in $d$, i.e., $\Theta(d)$  [^1] The notation $g(x)=\Theta(f(x))$ indicates that two functions $f(x)$ and $g(x)$ have the same asymptotic scaling with $x$, or more precisely, that there exists some constants $c_{1}$ and $c_{2}$ such that $c_{1}f(x)\leq g(x)\leq c_{2}f(x)$ for sufficiently large $x$. (see Fig. 1(a)). This is the case, for example, for the celebrated surface code [8, 9, 10], one of the leading candidates for practical FTQC due to its simple 2D layout and competitive error thresholds. Here, even the most basic task of fault-tolerantly preparing a logical $|{\overline{0}}\rangle$ state requires $\Theta(d)$ SE rounds [18], reducing the logical clock speed by a factor proportional to the code distance (typically 10 –100 [16, 14]). Alternative methods for FT state preparation exist [19], but incur a space-time trade-off and do not reduce the space-time volume. The same considerations also apply to state preparation and logical operations in many quantum low-density parity-check (QLDPC) codes [20, 21]. While there have been various efforts at addressing this challenge [5, 22], these alternative approaches introduce higher hardware complexity [23, 24, 21, 25] or necessitate certain properties of the underlying codes, such as the single shot QEC property [26, 27, 28, 29, 30, 31].
@@
Figure 1: Transversal algorithmic fault tolerance. (a) Conventional FT analysis separately examines each gadget (red boxes) in the circuit and ensures they are individually FT [4, 7, 32]. This typically requires $\Theta(d)$ syndrome extraction (SE) rounds to achieve FT. (b) Transversal algorithmic FT directly uses all accessible syndrome information up to a logical measurement (blue box), and guarantees FT of the logical measurement result, even if the gadgets are not individually FT and if future syndrome information is not yet accessible (partial decoding). We realize this through transversal operations, and only require a single SE round per logical operation.
@@
We introduce and develop an approach to FTQC that we refer to as “transversal algorithmic fault tolerance”, and show that it can lead to a substantial reduction in space-time cost for universal quantum computation. The key idea is to consider the fault tolerance of the algorithm as a whole (Fig. 1(b)) [33, 34, 35], allowing us to produce the correct classical logical measurement output with confidence in the presence of feed-forward operations, even though the intermediate quantum state is far from the ideal logical state. Contrary to the common assumption that fault-tolerant Pauli state preparation lower bounds the time cost of logical circuits by $d$ [10, 31, 36, 37], unless additional space-time trade-offs are made [19], we show that for any Calderbank-Shor-Steane (CSS) QLDPC code [6, 38], all transversal Clifford operations (including logical Pauli state preparation) can be performed with only constant time overhead per operation, provided that decoding can be implemented efficiently. We verify these results through a combination of proofs and circuit-level numerical simulations of our protocol, including a simulation of state distillation factories [39, 13], finding little change in error threshold compared to the standard surface code memory. Specializing to the surface code, our results reduce the number of SE rounds from $\Theta(d)$ to $\Theta(1)$, representing a direct reduction in the overall space-time overhead when classical decoding is sufficiently fast. This is particularly relevant for neutral atoms and trapped ions, where transversal gates have been natively implemented [40, 41, 42], and the operation timescales provide sufficient time to perform the more complex decoding problems.
Transversal Algorithmic Fault Tolerance
@@
Figure 2: Illustration of fault tolerance strategy. (a) Illustration of the measurement distribution remaining invariant under the application of Pauli initialization errors. (b) Logical quantum circuit with measurement and feed-forward. All logical operations are transversal and interleaved with a single SE round, instead of $d$ SE rounds. We must decode and commit mid-circuit to a measurement result for the top qubit, despite lacking complete syndrome information on the remaining qubits (partial decoding). (c) With the measurement result of the top qubit, a feed-forward operation is applied and the circuit continues execution. Upon reaching a new logical measurement, decoding is performed again on the whole circuit. The new decoding round may assign a different result to the top qubit, causing an inconsistency in feed-forward operations. (d) To restore consistency and reliably interpret the logical measurement results given past measurements, we apply a $\overline{Z}$ operator on the $|{\overline{0}}\rangle$ initial state, which acts trivially on $|{\overline{0}}\rangle$, but changes the interpreted logical measurement results to be consistent with before. This also leads to a re-interpretation of the new logical measurement result. Inset: Illustration of a single syndrome measurement error in the surface code, which causes a Pauli initialization error between distinct decoding rounds.
We focus on logical circuits comprised of Clifford operations and magic state inputs, where logical Clifford operations are transversal, i.e. implemented via depth-one quantum circuits [43, 44]. This is interleaved with SE rounds using ancilla qubits, which reveal error information on the data qubits and enable error correction. In addition to transversal gates [7], we refer to preparation of data qubits in $|0\rangle$ followed by one SE round as transversal state preparation, and $Z$ basis measurement of all data qubits as transversal measurement. To achieve universality, we allow teleporting in low-noise magic states with feed-forward operations based on past measurement results, and use the same Clifford operations above to prepare high quality magic states via magic state distillation [39]. We make use of CSS QLDPC codes, where each data or ancilla qubit interacts with a constant number of other qubits, thereby limiting error propagation, and each stabilizer generator consists of all $X$ or all $Z$ operators [38, 6]. Within this setting, our key result can be formulated as the following theorem:
Theorem 1 (informal): Exponential error suppression for transversal Clifford operations with a constant number of SE rounds per operation. For a transversal Clifford circuit with low-noise magic state inputs and feed-forward operations, that can be implemented with a given CSS QLDPC code family $\mathcal{Q}_{d}$ of growing code distance $d$, there exists a nonzero threshold $p_{\mathrm{th}}$, such that if the physical error rate $p<p_{\mathrm{th}}$ under the basic model of fault tolerance [5], then our protocol can perform logical operations with only a single SE round per operation, while suppressing the total logical error rate as $P_{L}=\exp(-\Theta(d))$.
The formal theorem statement and the corresponding proof can be found in Supplementary Materials [45]. Following the model in Ref. [5], our proof assumes the most likely error (MLE) decoder and fast classical computation. While this is not efficient for general hypergraph decoding problems, our simulations show that a threshold still exists for heuristic polynomial time decoders [45]. Fast decoders are an important area of future research, and we outline possible approaches towards this for practically-relevant circuits in the Methods section. We consider the local stochastic noise model, where we apply depolarizing errors on each data qubit every SE round and errors on each SE measurement, with a probability that decays exponentially in the weight of the total error event. Finally, we assume that all code patches are identical. While this theorem assumes low-noise magic state inputs, we conjecture that for the surface code, where methods for state injection are well-developed [46, 47, 48], the same approach can be applied to certain magic state distillation procedures [39] to reduce the space-time cost by a factor of $\Theta(d)$.
[…]
If one performs only a single SE round, as our approach prescribes, then a single measurement error on the randomly-initialized $X$ stabilizers can lead to a long, erroneous chain of $Z$ corrections (inset of Fig. 2(d)), bringing us far from the desired logical state. Indeed, with only $O(1)$ SE rounds and applying corrections, the resulting state is a mixture of product states and not close to any logical codeword [18]. With this unreliable state preparation, we may be concerned about incorrect logical measurements and feed-forward gates [49], which are crucial for achieving universality via magic state teleportation. Indeed, as we numerically demonstrate in Fig. 3(b) and show with an explicit example in SI, new syndrome information as the circuit continues execution has a high chance of changing our assignment of a logical measurement result (Fig. 2(b,c)), seemingly leading to an inconsistent Clifford feed-forward operation (Fig. 2(c)). For this reason, prior work analyzing circuits with transversal gates assumed that at least $d$ SE rounds separated Pauli state initialization and measurements or out-going qubits [31, 36, 37]. As we discuss in Methods, Pauli state initialization is used in many important algorithmic subroutines, and the $d$ SE rounds therefore lead to a substantial cost. Our approach overcomes this lower bound by instead considering the algorithm as a whole and exploiting the structure of initialization errors to fix these apparent inconsistencies in software.
The result of a quantum computation is fully characterized by its classical logical measurement distribution $P(z_{j},z_{j-1},...,z_{1})$. Our crucial observation is that errors which do not change this distribution, such as the initialization errors described above, do not affect the output (Fig. 2(a)), even if they cause the intermediate quantum state to be far from an ideal logical state.
The core reason for this is that universal quantum computation is realized here by an adaptive Clifford circuit on Pauli and magic state inputs. As the computation proceeds, Clifford feedforward gates are applied, such that the circuit executed so far is a fixed transversal Clifford circuit. A Clifford circuit maps Paulis to Paulis, thereby mapping both Pauli initialization errors and logical Pauli stabilizers identically to other Paulis (Fig. 2(d)). For each new measurement, if there is a product of measurements involving the new one that commutes with all Pauli initialization errors, then these initialization errors will not affect the product, and we can reliably and consistently infer the new measurement outcome [50, 51]. If no such product exists, a Pauli error that anti-commutes with a measurement can flip its result, naively causing a logical flip. However, the distribution is unchanged, because the anti-commutation of the measurement and corresponding logical stabilizer guarantees the result to already be 50/50 random. At an intuitive level, the reason this always works is because an $\bar{X}$ error does not do anything on a $|{\overline{+}}\rangle$ state.
We now provide additional elaboration of this result and show how it is concretely implemented. We consider the joint measurement distribution of the first $j$ logical measurements (Fig. 2(c)):
[…]
By considering the algorithm as a whole and leveraging the deterministic propagation of errors through transversal Clifford circuits, one can use the surrounding syndrome history to correct for noisy measurements (Fig. 1(b)) [15]. Because the resulting decoding hypergraph (characterizing how errors triggers syndrome checks) has bounded degree, the number of connected error clusters with size $s=\Theta(d)$ is bounded [5, 52], limiting the probability of a logical error caused by errors that do not intersect with any state initialization. However, we still need to account for error clusters that reach a state initialization.
More specifically, upon decoding with new syndrome information, there is a non-negligible chance that we change our assignment of the logical measurement result (green curves in Fig. 3(b)), seemingly leading to an inconsistent feed-forward (Fig. 2(c)). Crucially, however, the logical errors that lead to these inconsistencies are structured: since magic state inputs are prepared with low-noise, they do not lead to such inconsistencies. Therefore, the inconsistencies always correspond to a logical Pauli XX/$Z$ operator acting on the $|{\overline{+}}\rangle$$|{\overline{0}}\rangle$ initial state. Propagating this Pauli initialization error through the executed Clifford circuit results in a Pauli product.
There are two possibilities for how this error affects a logical measurement (Fig. 2(d)). First, if the measurement basis commutes with the resulting Pauli error (green), then it is unaffected and the result will be correct. Second, if the measurement basis anti-commutes with the resulting Pauli error (red), then the result will be flipped by the error. However, because correlated decoding ensures the correlations between current and prior measurements are correct, there is always a logical stabilizer associated with the $|{\overline{+}}\rangle$$|{\overline{0}}\rangle$ initial state, which flips the logical measurements identically as the initialization error. Applying this logical stabilizer when interpreting the logical measurement allows us to restore consistency with our previously-committed result (Fig. 2(d)). This process can be efficiently determined by solving a linear system of equations [45]. Failure to restore consistency requires an error cluster of size $d$ that is not connected to any initialization [45], which has a probability exponentially small in $d$. The above arguments show that both conditions for Eq. (1) to reproduce the ideal distribution are satisfied with exponentially-vanishing failure probability, proving Thm. 1 (see SI [45] for further details).
@@
Specializing to the surface code and utilizing the full transversal Clifford gate set accessible to the surface code (Methods), an immediate corollary of our main theorem is a threshold result for performing constant time logical operations with an arbitrary transversal Clifford circuit, when classical decoding is sufficiently fast. This result supports universal quantum computing when we allow magic state inputs prepared with sufficiently low noise.
Preparing high quality magic state inputs, in turn, can be performed simply with the same Clifford operations and easy-to-prepare non-fault-tolerant magic states [46, 47, 48], a procedure known as magic state distillation [39]. It will be interesting to extend our analysis to magic state distillation, where the ability to remove $d$ SE rounds from Pauli state initialization will be crucial to achieve the full savings [37, 31, 36]. The distillation factory and main computation can then be combined by applying our decoding approach to the joint system.
@@
Figure 3: Numerical verification of fault tolerance. (a) Simulation of circuit with repeated $\overline{ZZ}$ measurement (inset), where we commit mid-circuit to each measurement result of the logical ancilla using only the syndrome information up to that point. The total logical error rate as a function of circuit-level physical error rate $p$, for varying code distance $d$, shows clear threshold behavior. (b) Probability of inconsistent assignments of logical measurement results with and without the second step of our decoding strategy, as a function of code distance and for different physical error rates, for the same circuit as (a). Only with both steps do we observe exponential suppression of the logical error rate. (c) Comparison of two different methods for logical state preparation between three rotated surface codes and subsequent teleportation, for fixed circuit noise $p=0.3\%$. We use transversal gates (left) and lattice surgery (right), in both cases with only a single SE round. (d) With transversal gates, the error rate decreases exponentially with the code distance. With a single round of lattice surgery, the error rate instead increases linearly with code distance, as a single stabilizer measurement error affects the logical $\overline{ZZ}$ measurement result.
@@
Figure 4: $|{\overline{S}}\rangle$ state distillation factory. (a) Illustration of a $|{\overline{S}}\rangle$ state distillation factory based on the [[7,1,3]] Steane code, consisting of state initialization, layers of transversal CNOTs, followed by a teleported $\overline{S}$ gate. Each operation involves only a single SE round. Two of the CNOTs in the first layer act trivially and can be omitted. (b) The $|{\overline{S}}\rangle$ resource state is prepared via state injection at the first level, and via the first-level factory for the second level. (c) 1-level factory output state infidelity as a function of input state infidelity, for fixed circuit noise $p=0.1\%$ and varying levels of artificially injected $Z$ errors. The ideal curve is calculated assuming the gate operations in the factory are perfect. (d) Performance for one and two rounds of distillation, showing good agreement with the expected scaling.
@@
We first consider a simple circuit with intermediate logical measurements (inset of Fig. 3(a)) to evaluate the threshold and logical performance under partial syndrome information and without fault-tolerant state preparation. In this example, two logical qubits are transversally initialized in $|{\overline{+}}\rangle$, and an ancilla logical qubit is used to measure the $\overline{Z}\overline{Z}$ correlation a total of eight times, before the two logical qubits are transversally measured in the $\overline{Z}$ basis. While individual logical measurement results are random, a correct realization of this circuit should yield the same result for $\overline{Z}\overline{Z}$ each time, despite only using syndrome information accessible up to that point. We use the rotated surface code, a circuit-level depolarizing noise model [53, 15], a MLE decoder based on integer programming which is not efficient in general [54, 15], and employ the two-step process described above (see SI [45]).
@@
Figure 3(a-b) show the results of numerical simulations. We find that the total logical error rate shows characteristic threshold behavior, with an estimated threshold $\gtrsim 0.85\%$. This is close to the circuit-noise memory threshold, which can be understood from the fact that an SE round involves four times as many CNOT layers as a transversal CNOT, and therefore dominates the effective error rate. The number of SE rounds can be further optimized. For example, in Ref. [15], performing one SE round every four gate layers minimized the space-time cost per CNOT, suggesting that the practical improvement may be $\gtrsim 2d$ in some regimes  [^2] Four transversal gates and one SE round require 8 CNOT gates in total, compared to $16d+4$ with $d$ SE rounds.. In Fig. 3(b), we further compare the scaling of heralded failure rates in the presence and absence of the second step of our decoding procedure that restores consistency, as a function of code distance $d$. Here, the heralded failure rate is the probability that distinct rounds of decoding assign different values to the same logical measurement. We find that this additional step is necessary to achieve exponential suppression with the code distance in our setting where the logical Pauli states are prepared with only a single SE round.
@@
We now contrast our approach with lattice surgery in a single SE round setting [11, 12, 56, 57]. We consider the logical circuit in Fig. 3(c), where a GHZ state preparation circuit is followed by teleportation of the GHZ state to another set of logical qubits, and then measurement in the $\overline{Z}$ basis [45]. Using transversal gates with only a single SE round during $|{\overline{+}}\rangle$ and $|{\overline{0}}\rangle$ state preparation, and decoding each logical measurement with only accessible information at that stage, we find that the logical error rate decreases exponentially with the code distance, consistent with our FT analysis. In contrast, state preparation based on a single round of lattice surgery [56], which involves performing syndrome extraction with a larger code patch and then splitting it into three individual logical qubits, does not yield improved logical error rate as the code distance increases, as a single error can lead to incorrect inference of the $\overline{Z}\overline{Z}$ correlation of the GHZ state (SI). Unlike transversal measurements, logical information here is contained in noisy stabilizer products, which require repetition to reliably infer.